Gary runs on a path along the edge of a
rectangular park. The park is 1,560 feet
long and 750 feet wide. Gary runs the path
4 times every morning. How many miles
does Gary run? (1 mile = 5,280 feet)

Answer by nerdybill(7384)

Gary runs on a path along the edge of a
rectangular park. The park is 1,560 feet
long and 750 feet wide. Gary runs the path
4 times every morning. How many miles
does Gary run? (1 mile = 5,280 feet)
.
Perimeter = 2(1560+750) = 2(2310) = 4620 feet
.
Four times around:
4 * 4620 = 18480 feet
.
Miles ran:
18480/5280 = 3.5 miles